Write an equation of the line that passes through the origin and A(4,-5)
Answer by LinnW(1048)

The slope of the line through (0,0) and (4,-5) = (-5 -0)/(4 -0) = -5/4
So using m = -5/4 in y=mx+b, we know b = 0 since it is the intercept
we have y = (-5/4)x + 0 =
y = (-5/4)x